ennings Advertising Inc. reported the following in its December 31, 2018, balance sheet:

Equipment $ 600,000
Less: Accumulated depreciation—equipment $ 95,000

    
In a disclosure note, Jennings indicates that it uses straight-line depreciation over 12 years and estimates salvage value at 5% of cost. What is the average age of the equipment owned by Jennings?

Multiple Choice

  • 3.2 years.

  • 2 years.

  • 8.8 years.

  • 10 years.
Answer #1

Salvage value = 600,000*5% = 30,000

Depreciation expense per year = (Cost - Salvage value) /useful life

= (600,000-30,000)/12 = 47,500

Average age = Accumulated Depreciation/Depreciation expense per year

= 95,000/47,500

= 2 years